OpenStudy (gabylovesyou):

Evaluate: [(40 + 5) − 3^2] ÷ 9 • 2

OpenStudy (gabylovesyou):

8.6 43 8 2

OpenStudy (anonymous):

2

OpenStudy (gabylovesyou):

that was fast lol

OpenStudy (anonymous):

Thanks, let me explain it :P

OpenStudy (gabylovesyou):

k yes plis

OpenStudy (theviper):

First solve brackets:- [(40+5)-3^2]/9*2 [45-9]/9*2 36/9*2 4*2=8

OpenStudy (gabylovesyou):

so is it 8 or 2???????????

OpenStudy (anonymous):

(40+5) - 3^2 45 - 3x3 45 - 9/ 36 and 9x2 is 18/ 36/18/ 2

OpenStudy (theviper):

@AmauryP u made a mistake

OpenStudy (anonymous):

How?

OpenStudy (theviper):

It's\[\frac{36}{9}\times2\]

OpenStudy (theviper):

So it wold be:- \[\frac{72}{9}=8\]

OpenStudy (theviper):

u multiplied by the denomianator :P

OpenStudy (gabylovesyou):